OpenStudy (anonymous):

Write an equation of a line in point-slope form that passes through the point (1,2) and whose slope is 3.

OpenStudy (anonymous):

Yes. thats a good one. So do you know the equation of line if (x1,y1) = (1,2) and slope M = 3 ?

OpenStudy (anonymous):

i mean the formula used to find it.

OpenStudy (anonymous):

y=f(x) ???

jimthompson5910 (jim_thompson5910):

point slope form is y-y1 = m(x-x1) in this case m = 3 (slope) x1 = 1 (x coordinate of given point) y1 = 2 (y coordinate of given point